# Inequality in a model of capitalist economy

### v1

@@ Abstract
We analyze inequality aspects of the agent-based model of capitalist economy named Social Architecture of Capitalism that has been introduced by Ian Wright. The model contemplates two main types of agents, workers and capitalists, which can also be unemployed. Starting from a state where all agents are unemployed and possess the same initial wealth, the system, governed by a few simple rules, quickly self-organizes into two classes. After a transient, the model reproduces the statistics of many relevant macroeconomic quantities of real economies worldwide, notably the Boltzmann-Pareto regimes of the distributions of wealth and income. We perform extensive simulations testing the role of the model parameters (number of agents, total wealth, and salary range) on the resulting distribution of wealth and income, the social distribution of agents, and other stylized facts of the dynamics. Our main finding is that, according to the model, in an economy where total wealth is conserved and with a fixed average wage, the increase in wealth per capita comes with more inequality.
@@ I Introduction
The search of a first-principle model capable of elucidating the emergence and robustness of Pareto’s law [1] in various contexts is a central challenge in the field of Econophysics. Such a model would provide a fundamental understanding of the mechanisms that drive these distributions, transcending mere empirical observations to reveal the possible underlying principles.
@@
In his seminal work “The Social Architecture of Capitalism” [2], Ian Wright explores the intricate dynamics that results in the self-organization of a capitalist society with a given distribution of wealth among its members. Using an agent-based model known as Social Architecture (SA) model, Wright provides a theoretical framework aiming to explain the ubiquitous presence of Pareto’s law, which reflects that a small percentage of a population controls a large portion of a particular resource such as wealth or income. Wright’s work offers insights into how individual interactions and institutional structures collectively shape economic distributions. The model effectively reproduces the 80/20 rule  [^1] Very often regarded as the $80/20$ rule, meaning that 20% of the population owns 80% of the resources; the exact value of the ratio varies among countries or societies., and similar relations, which are key aspects of Pareto’s law, but, most remarkably, it naturally generates the two-pattern shape (Gamma distribution for the low wealth/income, power-law for the high wealth/income) of the wealth or income distributions, as originally observed by Pareto.
@@
Wright compares his results with real data, demonstrating that the model captures a significant portion of the dynamics of the capitalist system. In addition to the two regimes of wealth and income distributions, the model can reproduce the statistics of other economic indicators, such as firm size (power law), firm size growth rate (double exponential), firm demises per month (lognormal), and recession duration (exponential). These patterns emerge naturally and robustly from the simple set of rules that define the model. Hence, the SA model lays the foundation for exploring various economic phenomena, starting from a microscopic (or agent-based) dynamics controlled by a few parameters. Several follow-up studies have been published [4, 5, 6, 7, 8], but the SA model itself has not been thoroughly analyzed.
@@ II Model
The model implemented in this paper is the original SA proposed by Ian Wright [2]. However, we have chosen a slightly different notation and terminology to align with those commonly used in econophysics and agent-based modeling. The society consists of $N$ agents (labeled $i=1,\dots,N$ [^2] An agent is not necessarily an individual, but can represent other economic entities., where each agent $i$ has a positive integer quantity $w_{i}(t)$ representing wealth, which varies over time  [^3] While this quantity does not need to be an integer, we follow the original proposal, as this distinction is not critical to our analysis.. The population size $N$ and the total wealth of the system $W=\sum_{i=1}^{N}w_{i}(t)$ are conserved, so the wealth per capita $\overline{w}=W/N$ is constant. Agents can be in one of three classes: employees (working class), employers (capitalist class), or unemployed. Then, each agent is characterized by an index $e_{i}\neq i$, which identifies their employer: $e_{i}=j$ if agent $j$ is the employer of agent $i$, and $e_{i}=0$ if agent $i$ is unemployed or employer. Therefore, at any time $t$, the state of the entire economy is defined by the set of pairs $S(t)=\left\{\left(w_{i}(t),e_{i}(t)\right):1\leq i\leq N\right\}$. A firm consists of a set of employees and an employer, which is the only owner of the firm.
In addition to the system size $N$ and the wealth per capita $\overline{w}$, there are two other parameters: the minimum and maximum wages, $p_{a}$ and $p_{b}$, that each employee can receive. These four are the only parameters of the model. Random selections (of agents, wealth aliquots, wages, etc.) are made uniformly, unless otherwise specified.
@@
      If $w_{j}\geq\overline{p}$ (average of the distribution from which the wage (salary) is drawn), then: the agent $i$ is hired by $j$. Hence $j$ becomes an employer if it was previously unemployed.
@@
    Expenditure (on goods and services) – A random integer $w$ in the range $\left[0,w_{k}\right]$ is selected for a random agent $k\neq i$. Then, the following transfer to the market value $V$ occurs:
    $w_{k}\to w_{k}-w,\quad V\to V+w.$ (1)
@@
    Market revenue (from sales of goods and services) – Only if $i$ is not unemployed, the following transfer from the market value occurs:
@@
    $\begin{cases}w_{i}\to w_{i}+w,&\mbox{if $i$ is employer,}\\ w_{e_{i}}\to w_{e_{i}}+w,&\mbox{if $i$ is employee,}\end{cases}$
    where $w$ is a random integer number $w\in\left[0,V\right]$. In all cases, the quantity $w$ is counted as firm revenue $r_{i}$.
@@
      The number $m$ of employees to be fired is defined according to the formula $m=max(n_{i}-\frac{w_{i}}{\overline{p}},0)$, where $n_{i}$ is the number of employees of agent $i$.
@@
      Furthermore, if all the workers are fired, then the employer $i$ becomes an unemployed.
@@
The update is asynchronous, that is, the system state is updated each time each rule is applied. At the end of $N$ iterations of these six rules, $t$ is incremented by one unit (i.e., one month). Unless otherwise stated, the values of the parameters used in the simulations of the model are: system size $N=1000$ agents, wealth per captia $\overline{w}=100$, salary range $[p_{a},p_{b}]=[10,90]$. Typically, the maximum simulation time is $t_{max}=1000$ years, and the values of all state variables are recorded at the end of each year.
@@
III Preliminary analysis
The distribution of the population among the three classes of agents, along with the wealth distribution within each group, which emerges from the dynamics of the model, is displayed in Fig. 1. We observe a significant concentration of wealth within the capitalist class; however, it does not exhibit the typical pattern of wealth inequality —where most agents have no wealth— commonly found in binary exchange models from the existing literature when there is no intervention (such as taxes or assistance programs) [11, 12, 13, 14].
Figure 2: Distribution of wealth $f(w)$, (a) separated by agents’ type in a system of size $N=3200$, and (b) within the class of capitalists and (c) others, for two different population sizes ($N=100$ and $N=3200$). In all cases, $\overline{w}=100$, then for $N=100$ (3200), the total wealth in the system is $10^{4}$ ($3.2\times 10^{5}$). In the inset of (b), we show a detail of the tails in log-log scale. The segment with slope -2 is drawn for comparison. Data were accumulated over 100 realizations of 1000 years.
The cumulative distribution represented in Fig. 1 can be seen from a different perspective, as given by the probability density functions displayed in Fig. 2(a), which shows that, as an outcome of the model, employees and employers reside in separate wealth ranges. In Fig. 2(b), we focus on the distribution within the capitalist class, for two different system sizes. In both cases, the left-skewed shape is qualitatively similar, as well as the tails (shown in the inset). However, we observe a shift in the mode of the distribution to larger values when the number of agents increases, despite the fact that wealth per capita remains the same ($\overline{w}=100$). For both sizes of the system, the mode is well above $\overline{w}$, indicating wealth concentration. Moreover, increasing system size, which is accompanied by a greater injection of wealth into the system, enhances wealth concentration. Concomitantly, the distribution of wealth within the employee’s class (shown in Fig. 2(c)) suffers opposite impact, such that the mode decreases, and the peak narrows, meaning that wealth becomes more homogeneously distributed in the employee’s class.
In the following sections we will take a closer look on this impact of system size and wealth per capita on the economy dynamics. We will also see that a key parameter of the model is $R\equiv\overline{w}/\overline{p}$. In fact, first note that, according to the rules, the dynamics will not be impacted if wealth is scaled by an arbitrary quantity, for example, by $\overline{p}$, which only alters the monetary unit. In most cases, we will set $\overline{p}=100$. Given the chosen initial state, where agents start as unemployed with the same wealth $\overline{w}$, then, $R\geq 1$ is necessary to comply with $w_{j}>\overline{p}$ (rule 2(b)), to start hiring, and to prevent the system from becoming trapped in its initial state.
@@ IV.1 Influence of system size $N$
Figure 4: (a) Complementary cumulative distribution of wealth $F(w)$ of the full population and (b) corresponding Lorenz curves, for different system sizes $N$ indicated in the legend, with $\overline{w}=100$ and wage range $[10,90]$. In (a), the line with slope -1 is drawn for comparison. In (b), the inset shows the average annual (over 1000 years) Gini coefficient vs $N$.
In Fig. 3(a), we plot the wealth per capita across different classes as a function of $N$. The overall average wealth (indicated by symbols) shows a weak dependence on $N$. However, the average wealth of capitalists and the unemployed steadily increases with $N$, while that of employees decreases slightly. This trend suggests that as the population grows, capitalists accumulate more wealth and workers become more impoverished, although these effects stabilize at nearly constant levels for large systems. The concurrent increase in average wealth among the unemployed may be linked to firm crises.
Examining the square root of the variance of the wealth distribution (shown by error bars in Fig. 3), we find that it decreases with $N$ for employed and unemployed agents, indicating increased homogeneity within these groups. Conversely, the variance among capitalists rises with $N$, likely due to the extended asymptotic tail seen in Fig. 2, which suggests a divergent second moment. As $N$ increases keeping $\overline{w}$ constant (thus raising total wealth $W$), the range of capitalist wealth broadens, resulting in greater heterogeneity and an increased likelihood of encountering extremely wealthy agents, as illustrated in Fig. 4(a). Notice the change in the asymptotic behavior of $F(w)$ as $N$ increases, where the tail decays more slowly, approaching the law $F(w)\sim 1/w$, indicating a greater concentration of large fortunes within an oligarchic group. This trend is reflected in the Gini index, which increases gradually with $N$ and tends to saturation, as shown in the inset of Fig. 4(b).
@@ IV.2 Influence of the wealth per capita $\overline{w}$
Figure 3(b) shows how the distribution of wealth across the different classes (scaled by $\overline{w}$) is affected when changing the wealth per capita $\overline{w}$. The increase of $\overline{w}$ has a strong effect on the distribution of wealth, notability enhancing the capitalists’ wealth while reducing the wealth in the working and unemployed classes.
Figure 5: (a) Complementary cumulative distribution of wealth $F(w)$ of the full population and (b) corresponding Lorenz curves, for different levels of the wealth per capita $\overline{w}$, indicated in the legend. The population size is $N=1000$ and the wage range $[10,90]$. In (b), the inset shows the average annual Gini coefficient vs $\overline{w}$.
Upon closer examination of the cumulative distribution in Fig. 5(a), it is clear that the tail decays more slowly as $\overline{w}$ increases. Recalling that this second regime is associated with the capitalist class, according to Figs. 1 and 2, the observed flattening implies that capitalist wealth tends to concentrate in a more exclusive oligarchy with the higher $\overline{w}$. In other words, the probability of finding agents with very high wealth increases. This is an expected effect, also captured by the Lorenz curves in Fig. 5(b), showing that an increase in total wealth leads to a noticeable increase of the Gini coefficient, i.e., the inequality increases. This illustrates the common mistake of using a high gross domestic product as an indicator of progress and equality of opportunities!
For sufficiently large values of $\overline{w}$, (e.g., for $\overline{w}\gtrsim 300$ in the figure), there is a drastic change in the profile of $F(w)$, in which the second regime flattens. This implies the appearance of a disconnected very wealthy class. Moreover, let us remark that, in the extreme of $R\equiv\overline{w}/\overline{p}\gg 1$, the system can freeze with the condensation of the workforce in an oligarchy that concentrates most part of the wealth of the system and their firms cannot bankrupt, which is what happens in the case of the largest value of $\overline{w}$ plotted.

### v2

@@ Abstract
We analyze inequality aspects of the agent-based model of capitalist economy named Social Architecture of Capitalism that has been introduced by Ian Wright. The model contemplates two main types of agents, workers and capitalists, which can also be unemployed. Starting from a state where all agents are unemployed and possess the same initial wealth, the system, governed by a few simple rules, quickly self-organizes into two classes. After a transient, the model reproduces the statistics of many relevant macroeconomic quantities of real economies worldwide, notably the two regimes of the distributions of wealth and income. We perform extensive simulations testing the role of the model parameters (number of agents, total wealth, and salary range) on the resulting distribution of wealth and income, the social distribution of agents, and other stylized facts of the dynamics. Our main finding is that, according to the model, in an economy where total wealth is conserved and with a fixed average wage, the increase in wealth per capita comes with more inequality.
@@ I Introduction
The search of a first-principle model capable of elucidating the emergence and robustness of Pareto’s law Pareto (1896) is a central challenge in the field of Econophysics. Actually, as originally reported by Pareto, income and wealth real-world present two regimes: the upper tail follows a power-law (Pareto’s law), while the lower values display exponential behavior  Drăgulescu and Yakovenko (2001); Yakovenko (2024). Such first-principle model would provide a fundamental understanding of the mechanisms governing these properties, transcending mere empirical observations to reveal the possible underlying principles that drive economic inequality and its evolution over time. In this regard, agent-based models Chakraborti et al. (2011), particularly kinetic exchange models Aydiner et al. (2023), allow for the simulation of transactions among agents —such as individuals, firms, or other economic entities— that drive economic dynamics and help identify the underlying processes shaping these systems. This approach has paved the way for numerous efforts over the past few decades to explain a variety of phenomena, particularly the unequal distribution of wealth (see, for instance,  Boghosian (2014a); Aydiner et al. (2019); Chakraborti and Chakrabarti (2000); Dragulescu and Yakovenko (2000); Bouchaud and Mézard (2000) and Yakovenko and Rosser (2009); Chakraborti et al. (2011); Chakrabarti et al. (2013) for reviews). It is well known that unbiased random exchanges result in a condensed phase populated by the poorest agents Boghosian (2014a); Cardoso et al. (2023). To prevent such condensation and generate realistic distributions, extra mechanisms or regulatory controls have been introduced. These mechanisms can include the introduction of specific tax structures Lima et al. (2022); Calvelli and Curado (2023); Dias and Gonçalves (2024) or other wealth constraints, which, while effective in shaping outcomes, can make the results less purely emergent and more dependent on externally imposed rules. Let us give some examples. In random asset exchange models, the introduction of savings can mitigate condensation Chakraborti and Chakrabarti (2000), stabilizing the wealth distribution with a most probable value away from zero. The two-regime wealth/income distributions can be reproduced by the introduction of disorder in saving propensity Chatterjee et al. (2004), and also by wealth-based advantage mechanisms Calvelli and Curado (2023); Vallejos et al. (2018). When networks are considered instead of the more common mean-field setup (where all agents are potentially connected), topology can influence the wealth distribution Kohlrausch and Gonçalves (2024). Moreover, the power-law exponent can be altered by diverse factors, for instance by the distribution of saving propensities Aydiner et al. (2018), or by network position exchanges Aydiner et al. (2019). In the present paper, using agent based simulations, we investigate a model that, simultaneously gives rise to two key phenomena: the emergence of the dual-regime structure in wealth and income distributions, alongside the development of a two-class society.
@@
In his seminal work “The Social Architecture of Capitalism” Wright (2005), Ian Wright explores the intricate dynamics that results in the self-organization of a capitalist society with a given distribution of wealth among its members. Using an agent-based model known as Social Architecture (SA) model, Wright provides a theoretical framework aiming to explain the ubiquitous presence of Pareto’s law, which reflects that a small percentage of a population controls a large portion of a particular resource such as wealth or income. Wright’s work offers insights into how individual interactions and institutional structures collectively shape economic distributions. The model effectively reproduces the so called, “80/20 rule” 111 The $80/20$ rule derives its name from Pareto’s observation of the Italian economy, where he found that roughly 20% of the population owned $\sim 80$% of the resources. The term is used generically, although the ratio $f:(1-f)$ may vary across different countries and societies, reflecting a broader pattern in which a small proportion of inputs or entities accounts for a large share of outcomes or resources Boghosian (2014b, a)., but, most notably, it naturally generates the two-pattern shape of the wealth or income distributions (exponential-like, for low wealth/income, power-law for high wealth/income).
@@
Wright compares his results with real data, demonstrating that the model captures a significant portion of the dynamics of the capitalist system. In addition to the two regimes of wealth and income distributions, the model can reproduce the statistics of other economic indicators, such as firm size (power law), firm size growth rate (double exponential), firm demises per month (log-normal), and recession duration (exponential). These patterns emerge naturally and robustly from the simple set of rules that define the model. Hence, the SA model lays the foundation for exploring various economic phenomena, starting from microscopic (or agent-based) dynamics controlled by a few parameters. Several follow-up studies have been published Cottrell et al. (2009); Wright (2009); Lavička et al. (2010); Isaac (2019); Villarreal (2022); Greenberg and Gao (2024), but the SA model itself has not been thoroughly analyzed.
@@ II Model
The model implemented in this paper is the original SA proposed by Ian Wright Wright (2005). However, we have chosen a slightly different notation and terminology to align with those commonly used in econophysics and agent-based modeling. The society consists of $N$ agents (which are labeled $i=1,\dots,N$ and are not necessarily individuals, but can represent other economic entities), where each agent $i$ has a positive integer quantity $w_{i}(t)$ representing wealth, which varies over time. Here wealth is defined as an amount of money, assets or goods that can be eventually transformed into money, owned by each person in a population. Although this quantity does not need to be an integer, we follow the original proposal, as this distinction is not critical to our analysis. The size of the population, $N$, and the total wealth of the system, $W=\sum_{i=1}^{N}w_{i}(t)$, are conserved, so the wealth per capita $\overline{w}=W/N$ is constant. Agents can be in one of three classes: employees (working class), employers (capitalist class), or unemployed. Then, each agent is characterized by an index $e_{i}\neq i$, which identifies their employer: $e_{i}=j$ if agent $j$ is the employer of agent $i$, and $e_{i}=0$ if agent $i$ is unemployed or employer. Therefore, at any time $t$, the state of the entire economy is defined by the set of pairs $S(t)=\left\{\left(w_{i}(t),e_{i}(t)\right):1\leq i\leq N\right\}$. A firm consists of a set of employees and an employer, which is the only owner of the firm.
The model has four parameters: system size $N$, wealth per capita $\overline{w}$, and the minimum ($p_{a}$) and maximum ($p_{b}$) wages employees can receive. Unless stated otherwise, all random selections (agents, wealth aliquots, wages, etc.) are uniform.
Besides agents (employees and employers), the model includes a market value $V$, representing goods and services produced by firms. $V$ grows with consumer expenditures and decreases with firm revenues. Firms compete for a share of this pool rather than receiving direct payments.
@@
      If $w_{j}\geq\overline{p}$ (average wage, or salary, from the distribution of wages), then: the agent $i$ is hired by $j$. If $j$ was unemployed, it becomes hence an employer.
@@
    Expenditure (on goods and services) – A random integer $w$ in the range $\left[0,w_{k}\right]$ is selected for a random agent $k\neq i$. This amount, representing the agent’s expenditure, is subtracted from their wealth and added to the market value, so:
    $w_{k}$$\to$ $w_{k}-w,$
    $V$$\to$ $V+w.$
@@
    Market revenue (from sales of goods and services) – Only if $i$ is not unemployed, a random integer $w\in[0,V]$ is transferred from the market value to the agent $i$:
@@
    $\begin{cases}w_{i}\to w_{i}+w,&\mbox{if $i$ is an employer,}\\ w_{e_{i}}\to w_{e_{i}}+w,&\mbox{if $i$ is an employee.}\end{cases}$
    In all cases, $w$ counts as revenue for the firm owner, either if $i$ is the owner itself or if $i$ is an employee (in which case $e_{i}$ is the owner).
@@
      The number $m$ of employees to be fired is defined according to the formula $m=max(n_{i}-\frac{w_{i}}{\overline{p}},0)$, where $n_{i}$ is the number of agents employed by agent $i$.
@@
      Furthermore, if all workers are fired, then the employer $i$ becomesunemployed.
@@
The update is asynchronous, that is, the system state is updated each time each rule is applied. At the end of $N$ iterations of these six rules, $t$ is incremented by one unit (i.e., one month). Unless otherwise stated, the values of the parameters used in the simulations of the model are: system size $N=1000$ agents, wealth per capita $\overline{w}=100$, salary range $[p_{a},p_{b}]=[10,90]$. Typically, the maximum simulation time is $t_{max}=1000$ years, and the values of all state variables are recorded at the end of each year. For statistical calculations, data were recorded after reaching a steady state, discarding a short transient period.
III Preliminary analysis
The distribution of the population among the three classes of agents, along with the wealth distribution within each group, which emerges from the dynamics of the model, is displayed in Fig. 1. We observe a significant concentration of wealth within the capitalist class; however, it does not exhibit the typical pattern of wealth inequality —where most agents have no wealth— commonly found in binary exchange models from the existing literature when there is no intervention (such as taxes or assistance programs) Moukarzel et al. (2007); Boghosian (2014b); Cardoso et al. (2020); Lima et al. (2022).
@@ III Preliminary analysis
Figure 2: Complementary cumulative distributions of the wealth per capita within each class. We used $N=102400$, $\overline{w}=100$, and salary range $[10,90]$. The segment with slope -1, which represents an upper bound, is drawn for comparison. Data of a single realization were accumulated over 1000 years.
The separate distributions of wealth per capita within is class is shown in Fig. 2, exhibiting a long tail for the capitalist class, with workers and unemployed present an abrupt cut-off. This decomposition helps to understand the origins of the two regimes observed in the global distribution of wealth, as shown in Fig. 1.
In the following sections we will take a closer look on this impact of system size and wealth per capita on the economy dynamics. We will also see that a key parameter of the model is $R\equiv\overline{w}/\overline{p}$. In fact, first note that, according to the rules, the dynamics will not be impacted if wealth is scaled by an arbitrary quantity, for example, by $\overline{p}$, which only alters the monetary unit. Typically, we will use $\overline{w}=100$ and $\overline{p}=50$, unless indicated differently. Given the chosen initial state, where agents start as unemployed with the same wealth $\overline{w}$, then, $R\geq 1$ is necessary to comply with $w_{j}\geq\overline{p}$ (rule 2(b)), to start hiring, and to prevent the system from becoming trapped in its initial state.
@@ IV.1 Influence of system size $N$
Examining the square root of the variance of the wealth distribution (shown by error bars in Fig. 3), we find that it decreases with $N$ for employed and unemployed agents, indicating increased homogeneity within these groups. Conversely, the variance among capitalists rises with $N$, likely due to the extended asymptotic tail seen in Fig. 2, which suggests a divergent second moment in the limit $N\to\infty$. As $N$ increases keeping $\overline{w}$ constant (thus raising total wealth $W$), the range of capitalist wealth broadens, resulting in greater heterogeneity and an increased likelihood of encountering extremely wealthy agents, as illustrated in Fig. 4(a). Notice the change in the asymptotic behavior of $F(w)$ as $N$ increases, where the tail decays more slowly, approaching the law $F(w)\sim 1/w$, indicating a greater concentration of large fortunes within an oligarchic group. This trend is reflected in the Gini index (or Gini coefficient), which increases gradually with $N$ and tends to saturation, as shown in the inset of Fig. 4(b). The same trend is reflected in the Kolkata index Banerjee et al. (2020) also shown in the inset.
Figure 4: (a) Complementary cumulative wealth distribution of the full population, $F(w)$, and (b) corresponding Lorenz curves, for different system sizes $N$ indicated in the legend, with $\overline{w}=100$ and wage range $[10,90]$. In (a), the line with slope -1 is drawn for comparison. In (b), the inset shows the average annual (over 1000 years) Gini and Kolkata indexes vs $N$. In the current and subsequent figures, each distribution was built with values recorded at the end of each year, accumulated over the years of a single realization. For the inequality indices, we present the average of the annual values (symbols) and their standard deviation (represented by bars).
@@ IV.2 Influence of the wealth per capita $\overline{w}$
Figure 3(b) shows how the distribution of wealth across the different classes (scaled by $\overline{w}$) is affected when changing the wealth per capita $\overline{w}$. The increase of $\overline{w}$ has a strong effect on the distribution of wealth, notability enhancing the relative wealth of the capitalists while reducing the relative wealth in the working and unemployed classes.
Upon closer examination of the cumulative distribution in Fig. 5(a), it is clear that the tail decays more slowly as $\overline{w}$ increases. Recalling that this second regime is associated with the capitalist class, according to Figs. 1 and 2, the observed flattening implies that capitalist wealth tends to concentrate in a more exclusive oligarchy with the higher $\overline{w}$. In other words, the probability of finding agents with very high wealth increases. This is an expected effect, also captured by the Lorenz curves in Fig. 5(b), showing that an increase in total wealth leads to a noticeable increase of the Gini index, i.e., the inequality increases. This illustrates the common mistake of using a high gross domestic product as an indicator of progress and equality of opportunities!
Figure 5: (a) Complementary cumulative wealth distribution of the full population, $F(w)$, and (b) corresponding Lorenz curves, for different levels of the wealth per capita $\overline{w}$, indicated in the legend. The population size is $N=1000$ and the wage range $[10,90]$. In (b), the inset shows the average annual Gini and Kolkata indexes vs $\overline{w}$.
For sufficiently large values of $\overline{w}$, (e.g., for $\overline{w}\gtrsim 300$ in the figure), there is a drastic change in the profile of $F(w)$, where the second regime flattens. This indicates the appearance of a very wealthy class, separated by a gap in the wealth distribution. Moreover, let us remark that, in the extreme of $R\equiv\overline{w}/\overline{p}\gg 1$, the system can freeze with the condensation of the workforce in an oligarchy that concentrates most part of the wealth of the system and their firms cannot bankrupt, which is what happens in the case of the largest value of $\overline{w}$ plotted.